OpenStudy (anonymous):

The cosine of 48°, to the nearest hundredth, is 0.67. To the nearest hundredeth, what is the cosine of 96°? How do find this out with out a calculator?

OpenStudy (asnaseer):

96 = 2*48 use formulae for cos(2x)

OpenStudy (anonymous):

yeah use the formula for cos(2x) \[ \cos 2x=2\cos ^{2}x-1\] x=48 this gives the value approx. to '-0.1391'